Bob invested $22,000, part at 18% and part at 14%. If the total interest at the end of the year is $3,360, how much did he invest at 18%?

Answer by checkley71(8403)

.18x+.14(22,000-x)=3,360
.18x+3,080-.14x=3360
.04x=3360-3080
.04x=280
x=280/.04
x=7,000 answer of amount invested @ 18%.
22,000-7,000=15,000 invested @ 14%.
proof
.18*7,000+.14*15,000=3,360
1,260+2,100=3,360
3,360=3,360